OpenStudy (anonymous):

Use seperation of variables (dy/dx) = (y/x) and y = 2 when x = 2

OpenStudy (amistre64):

feel free to use it then

OpenStudy (amistre64):

\[\int y^{-1}dy=\int x^{-1}dx\]

OpenStudy (anonymous):

my teacher went from e^lny = e^lnx+c to y = e^lnx*e^c can you ehy it is like that please

OpenStudy (amistre64):

ehy?

OpenStudy (anonymous):

explain why*

OpenStudy (amistre64):

you see that this thing ints up into lny = lnx + c right?

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

the inverse of ln IS e ... it means that they undo each other

OpenStudy (amistre64):

e^lny = y e^(lnx+c) is a bit trickier to notice but its based on exponent rules b^n * b^m = b^(n+m) e^(lnx+c) = e^lnx * e^c =x * e^c

OpenStudy (amistre64):

e^c is just some constant so they absorb it into one C y = x*C

OpenStudy (anonymous):

oh! I see it now ty

OpenStudy (amistre64):

yep